OpenStudy (anonymous):

Determine the slope of the line represented by the equation: 3x+4y=12

OpenStudy (anonymous):

is it -4/3

OpenStudy (rmrjr22):

y = mx + b m = slope

OpenStudy (rmrjr22):

y = 3-(3/4)x

OpenStudy (anonymous):

is it -4/3?

OpenStudy (rmrjr22):

-3/4

OpenStudy (rmrjr22):

you get y on one side in order to get the formual y = mx + B the way you did it was, you set x = 12 - 4y then you divided by 3

OpenStudy (rmrjr22):

Rate please if you can :)

OpenStudy (anonymous):